The substation bus bar is made up of made 1.5 inches round copper bars 30 ft long. What is the resistance of each bar?


1
Expert's answer
2022-03-01T20:46:01-0500

Radius of copper bar= 0.0381/2 = 0.019 meter

Area of copper bar = 3.4 (0.019)2= 0.059 m2

L= 9.144 m

The density of copper in SI =  8830 Kg /m3

Resistance of copper bar = Density of copper *(L/A)

Resistance of copper bar =8830 *(9.144/0.059)= 56.96 ohm